OpenStudy (zned6559):

A salt is known to be an alkali metal fluoride. A quick approximate freezing point determination indicates that 4 g of the salt dissolved in 100 g of water to produce a solution that freezes at about -1.4C. What is the identity of the salt. (Hint: Alkali fluoride salts are usually electrolytes)

OpenStudy (anonymous):

use the freezing point depression formula for this one: delta T = i * m * K where K is a constant, m is the molality (mol solute/kg solvent), and i is the van'hoff factor the van hoff factor is the number of ions that your salt dissociates into. Since it's an ALKALI flouride salt, how many ions? k is just a constant, you get it from a table in your textbook somewhere So you have everything to solve for the molality of the solution, once you did that, multiplying it by the mass of water to find the mols of the salt. Take the mass of the salt and divide by this mols to figure out the molar mass, and then compare it with the periodic table to identify the salt. \[\frac{ mole solute }{ kg solvent } \times mass of water \times = mol solute\] \[massofsolute \times \frac{ 1 }{ moleofsolute } = molarmass\]